# Non-Abelian anyons in a periodically-driven Abelian model

### 06 Feb 2024 v1

@@ I Quasiparticle mapping
$\hat{\bm{Z}}_{p}^{\psi}=-i\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime},\quad\hat{\bm{X}}_{p}^{\psi}=\left(\prod_{p^{\prime}\in\mathrm{JW}(p)}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\right)\hat{\gamma}_{p}^{\prime},$ (S12)
where JW$(p)$ is a Jordan-Wigner string with the convention shown in Fig. S1: the summation involves all plaquettes scanned left-to-right, row-by-row from the bottom of the lattice up, until $p$, thus following the path JW$(p)$ depicted in S1. In Eq. (S12), we have introduced the Majorana real-fermion operators
@@
where $\hat{f}_{p}^{\dagger}$ and $\hat{f}_{p}$ are creation and annihilation operators for a complex fermion ($\psi$). The two-spin terms of Eq. (S10d) then become
$\hat{X}_{p,R}\hat{Z}_{p,B}=i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+x}$ (S14a)
@@
$\hat{X}_{p,L}\hat{Z}_{p,T}=\left[\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\right]i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+y}.$ (S14b)
@@
For two-spin terms inducing fermion motion across the boundaries horizontally and vertically, respectively, the mapping yields
@@
$\hat{X}_{p,R}\hat{Z}_{p,B}=\left(\prod_{\begin{subarray}{c}v\in{\rm Ab}(p)\end{subarray}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\right)i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+x}\hat{\bm{Z}}_{x}^{W},$ (S15b)
@@
$\hat{X}_{p,L}\hat{Z}_{p,T}=\left(\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\right)i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+y}\hat{\bm{Z}}_{y}^{W},$ (S15c)
@@
involving the ‘Wilson-loop pseudospins’ $\hat{\bm{Z}}_{r}^{W}$, and where Ab$(p)$ indicates all vertices above the row of plaquettes to which $p$ belongs. In the absence of terms in the Hamiltonian that can induce a change of superselection sector (through tunnelling of a $e$-boson all around the torus), the choice of sector determined by the eigenvalues of the Wilson loop operators $\hat{W}_{x}$ and $\hat{W}_{y}$ then implies the choice of either periodic ($\bm{Z}_{x}^{W}=+1$, $\bm{Z}_{y}^{W}=+1$) of anti-periodic ($\bm{Z}_{x}^{W}=-1$, $\bm{Z}_{y}^{W}=-1$) boundary condition for the fermions along each direction. Defining the string of boson parities
$\hat{\mathcal{S}}_{p}^{(y)}=\left[\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\right],$ (S16)
@@
the driven toric-code Hamiltonian $\hat{H}_{\rm tc}+\hat{H}_{d}(t)$, with $\hat{H}_{\rm tc}$ of Eq. (1) and $\hat{H}_{d}(t)$ of Eq. (2), is then mapped, in the bulk, to
$\hat{H}(t)=$ $-\frac{g}{2}\sum_{v}(-1)^{\hat{b}^{\dagger}_{v}\hat{b}_{v}}$ (S17)
$+\sum_{p}\Big{\{}\frac{g}{2}\left[1+(-1)^{\hat{b}_{v(p)}^{\dagger}\hat{b}_{v(p)}}\right]i\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime}$
$+f_{x}(t)i\hat{\gamma}_{p}\hat{\gamma}_{p+x}^{\prime}+f_{y}(t)\hat{\mathcal{S}}_{p}^{(y)}\ i\hat{\gamma}_{p}\hat{\gamma}_{p+y}^{\prime}\Big{\}}$ (S18)
Eq. (S22) applies to bulk fermions: We did not explicitly include the corrections for the boundary terms, which we dump on the choice of boundary condition (we also assume that $e$ bosons are present only in the bulk). We next make the assumption that, whenever an $e$-boson pair is present at vertices $v$ and $v^{\prime}$, such that $\hat{b}_{v}^{\dagger}\hat{b}_{v}=\hat{b}_{v^{\prime}}^{\dagger}\hat{b}_{v^{\prime}}=1$, this corresponds to using a slightly modified toric-code Hamiltonian in Eq. (1) with $g\hat{A}_{v}\to-g\hat{A}_{v}$ and $g\hat{A}_{v^{\prime}}\to-g\hat{A}_{v^{\prime}}$, namely
$\hat{H}_{\mathrm{tc}}=\frac{g}{2}(\hat{A}_{v}+\hat{A}_{v^{\prime}})-\frac{g}{2}\sum_{v^{\prime\prime}\neq v,v^{\prime}}\hat{A}_{v^{\prime\prime}}\ -\frac{g}{2}\sum_{p}\hat{B}_{p},$ (S19)
such that the state containing the $e$ pair is the ground state with respect to the modified Hamiltonian. In this case, the Hamiltonian (S22), for a fixed configuration of $e$ bosons defined by a boson Fock state |$\ket{\vec{n}^{e}}=(\hat{b}^{\dagger}_{v_{1}})^{n_{1}}\dots(\hat{b}_{v_{N}}^{\dagger})^{n_{N}}\ket{0_{\rm tc}}$, gives the fermion Hamiltonian
$\hat{H}_{\psi}(t)=$ $\bra{\vec{n}^{e}}\hat{H}_{\rm tc}+\hat{H}_{d}(t)\ket{\vec{n}^{e}}$ (S20)
$=$ $\sum_{p}\Big{\{}gi\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime}+f_{x}(t)i\hat{\gamma}_{p}\hat{\gamma}_{p+x}^{\prime}$ (S21)
$\qquad+f_{y}(t)\hat{\mathcal{S}}_{p}^{(y)}\ i\hat{\gamma}_{p}\hat{\gamma}_{p+y}^{\prime}\Big{\}}$ (S22)
@@
Re-expressing the Majorana operators in terms of complex fermions $\hat{f}_{p}$ and $\hat{f}_{p}$ via Eq. (S13), defining the fermion chemical potential $\mu_{\psi}=g$ and neglecting constant energy shifts then gives the fermion Hamiltonian of Eq. (4) of the main text.
@@
Consider now $e$-boson (vortex) movement. This is induced by the single-spin operator $\hat{Z}_{i}$. In the quasiparticle mapping, for spins in the bulk, this is mapped to
$\hat{Z}_{v,R}=(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+x}^{\dagger}+\hat{b}_{v+x}),$
$\hat{Z}_{v,T}=\hat{\tilde{\mathcal{S}}}_{v}^{(y)}(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+y}^{\dagger}+\hat{b}_{v+y}),$ (S23)
with
@@ II Floquet engineering
In this section, the high-frequency effective Hamiltonian of Eq. (5), produced by the driven Hamiltonian $\hat{H}_{\psi}(t)$ of Eq. (3), is derived.
Following Floquet’s Theorem, the evolution operator generated at time $t$ by the time-periodic $\hat{H}_{\psi}(t)=\hat{H}_{\psi}(t+T)$, starting from initial time $t_{0}$, can be decomposed according to  [27]
$\hat{U}(t)=\hat{K}(t)e^{-i\hat{H}_{F}(t-t_{0})}\hat{K}^{\dagger}(t_{0}),$ (S25)
in terms of an effective time-independent Floquet Hamiltonian $\hat{H}_{F}$ and a time-periodic micromotion operator $\hat{K}(t)=\hat{K}(t+T)=e^{\hat{G}(t)}$, generated by $\hat{G}(t)$. Both $\hat{H}_{F}$ and $\hat{G}(t)$ can be determined order by order in a high-frequency expansion in powers of $\omega^{-n}$ [30, 25]. The leading order is given by
@@
$\hat{G}(t)=-\sum_{m\neq 0}\frac{e^{im\omega t}\hat{H}_{m}}{m\omega},$ (S26b)
where we have indicated the Fourier components of $\hat{H}_{\psi}(t)$ with
@@
To compute the expansion, we first represent the Hamiltonian $\hat{H}_{\psi}(t)$ in a rotating frame defined by the gauge transformation
$\hat{R}(t)=\exp\left(-it\frac{\omega}{2}\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}\right).$ (S28)
At the end of the period, $\hat{R}(T)$ reduces to the total fermion parity of the system, $\hat{R}(T)=\exp\left(-i\pi\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}\right)$. Since the latter is a conserved quantity, $R(t)$ commutes with the Floquet Hamiltonian, $[\hat{R}(T),\hat{H}_{F}]=0$. Moreover, it acts trivially within the physical subspace of even-parity states, and thus only contributes to their micromotion dynamics. The transformed Hamiltonian $\hat{H}^{\prime}_{\psi}(t)=\hat{R}^{\dagger}(t)\hat{H}_{\psi}(t)\hat{R}(t)-i\hat{R}^{\dagger}\partial_{t}\hat{R}(t)$ reads
@@
$+\!\!\sum_{r\in\{x,y\}}d_{r}(t)\hat{\mathcal{S}}_{p}^{(r)}\big{(}e^{i\omega t}\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{]}.$ (S29)
Inserting the drive $d_{r}(t)$ of Eq. (4) gives
$\hat{H}^{\prime}_{\psi}(t)=\sum_{p}\Big{\{}(\mu_{\psi}-\omega/2)\hat{f}_{p}^{\dagger}\hat{f}_{p}$
$-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}J\big{(}e^{i\omega t}\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}$
$-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\Delta\big{[}e^{i\phi_{r}}\big{(}\hat{f}_{p}\hat{f}_{p+r}+e^{i\omega t}(\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\hat{f}_{p+r}^{\dagger}\hat{f}_{p})$
$+e^{i2\omega t}\hat{f}_{p}\hat{f}_{p+r}\big{)}+\mathrm{H.c.}\big{]}\Big{\}}.$ (S30)
Computing the time average as per Eq. (S26a) yields the time independent part of Eq. (S30),
$\hat{H}_{0}=\sum_{p}\Big{\{}-\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\big{(}J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}\\ +\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{\}},$ (S31)
where we have defined the effective chemical potential $\mu=-(\mu_{\psi}-\pi/T)$. The kick operator $\hat{G}(t)$ to leading order is determined by Eq. (S26b), by $\hat{R}(t)$ and by the Fourier components of $\hat{H}^{\prime}_{\psi}(t)$, which from Eq. (S30) read

### 28 Feb 2024 v2

@@ I Quasiparticle mapping
$\hat{\bm{Z}}_{p}^{\psi}\leftrightarrow-i\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime},\quad\left[\prod_{p^{\prime}\in\mathrm{JW}(p)}\hat{\bm{Z}}_{p^{\prime}}^{\psi}\right]\hat{\bm{X}}_{p}^{\psi}\leftrightarrow\hat{\gamma}_{p}^{\prime},$ (S13)
where JW$(p)$ is a Jordan-Wigner path with the convention shown in Fig. S1(b) [and exemplified in Fig. S3(b)-(c) for vertical processes in the bulk]: the summation involves all plaquettes scanned left-to-right, row-by-row from the top-left corner of the lattice down, until $p$, thus following the path JW$(p)$ depicted in S1(b). In Eq. (S13), we have introduced the Majorana real-fermion operators
@@
which are Hermitian and satisfy $\hat{\gamma}_{p}^{2}=(\hat{\gamma}_{p}^{\prime})^{2}=1$ and, for $p\neq p^{\prime}$, $\{\hat{\gamma}_{p},\hat{\gamma}_{p^{\prime}}\}=\{\hat{\gamma}_{p}^{\prime},\hat{\gamma}_{p^{\prime}}^{\prime}\}=\{\hat{\gamma}_{p},\hat{\gamma}_{p^{\prime}}^{\prime}\}=0$. The operators $\hat{f}_{p}^{\dagger}$ and $\hat{f}_{p}$ are creation and annihilation operators for a complex fermion (representing $\psi$). From Eq. (S7), the toric-code Hamiltonian is then mapped to
$\hat{H}_{\rm tc}\leftrightarrow\frac{g}{2}\sum_{p}\left[1+(-1)^{\hat{b}_{v(p)}^{\dagger}\hat{b}_{v(p)}}\right]i\gamma_{p}\gamma_{p}^{\prime}.$ (S15)
In terms of complex-fermion operators $\hat{f}_{p}$ and $\hat{f}_{p}^{\dagger}$, Eq. (S15) can be expressed as
$\hat{H}_{\rm tc}\leftrightarrow 2g\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}+g\sum_{v}\hat{b}_{v}^{\dagger}\hat{b}_{v}\\ -2g\sum_{p}\hat{b}_{v(p)}^{\dagger}\hat{b}_{v(p)}\hat{f}_{p}^{\dagger}\hat{f}_{p},$ (S16)
as reported in the main text, where we also used the hardcore-boson constraint to express the boson parities as $(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}=1-2\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and we neglected constant energy shifts. The two-spin terms of Eq. (I) become
$\hat{X}_{p,R}\hat{Z}_{p,B}\leftrightarrow i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+x}$ (S17a)
@@
$\hat{X}_{v(p),R}\hat{Z}_{v(p),B}\leftrightarrow\left[\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\right]i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+y}.$ (S17b)
@@
The different stages of the mapping leading to Eq. (S17b) are sketched in S3 . For two-spin terms inducing fermion motion across the boundaries horizontally and vertically, respectively, the mapping yields
@@
$\hat{X}_{p,R}\hat{Z}_{p,B}\leftrightarrow\left[\prod_{\begin{subarray}{c}v^{\prime}\in{\rm Ab}(v)\end{subarray}}(-1)^{\hat{b}_{v^{\prime}}^{\dagger}\hat{b}_{v^{\prime}}}\right]i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+x}\hat{\bm{Z}}_{x}^{W},$ (S18a)
@@
$\hat{X}_{v(p),R}\hat{Z}_{v(p),B}\leftrightarrow\left[\prod_{v^{\prime}\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v^{\prime}}^{\dagger}\hat{b}_{v^{\prime}}}\right]i\hat{\gamma}_{p}\hat{\gamma}^{\prime}_{p+y}\hat{\bm{Z}}_{y}^{W},$ (S18b)
@@
involving the ‘Wilson-loop pseudospins’ $\hat{\bm{Z}}_{r}^{W}$, and where Ab$(v)$ indicates all vertices in the vertex rows above the row to which $v$ belongs. In the absence of terms in the Hamiltonian that can induce a change of superselection sector (through tunnelling of a $e$-boson on a full non-contractible loop around the torus), the choice of sector determined by the eigenvalues of the Wilson loop operators $\hat{W}_{x}$ and $\hat{W}_{y}$ then implies the choice of either periodic ($\bm{Z}_{x}^{W}=+1$, $\bm{Z}_{y}^{W}=+1$) of anti-periodic ($\bm{Z}_{x}^{W}=-1$, $\bm{Z}_{y}^{W}=-1$) boundary condition for the fermions along each direction. Defining the string of boson parities
$\hat{\mathcal{S}}_{p}^{(y)}=\left[\prod_{v^{\prime}\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v^{\prime}}^{\dagger}\hat{b}_{v^{\prime}}}\right],$ (S19)
@@
the driven toric-code Hamiltonian $\hat{H}(t)=\hat{H}_{\rm tc}+\hat{H}_{d}(t)$, with $\hat{H}_{\rm tc}$ of Eq. (1) and $\hat{H}_{d}(t)$ of Eq. (2), is then mapped, in the bulk, to
Figure S3: Sketch of the different stages of the quasiparticle mapping for vertical fermion processes in the bulk [Eqs. (S11d) and (S17b)]. A two-spin term $\hat{Z}_{v,B}\hat{X}_{v,R}$ inducing tunnelling and pairing of the toric-code fermionic quasiparticles on plaquettes $p$ and $p+y$ is first mapped to a product of pseudospin operators $\hat{\bm{Z}}_{p}^{\psi}$, $\hat{\bm{X}}_{p}^{\psi}\hat{\bm{X}}_{p^{\prime}}^{\psi}$ and $\hat{\bm{Z}}_{v}^{e}$ (b). Then, such products are mapped to products of hardcore-boson operators on vertices ($\hat{b}_{v}$ and $\hat{b}_{v}^{\dagger}$, describing $e$ bosons) and Majorana-fermion operators on plaquettes ($\hat{\gamma}_{p}$ and $\hat{\gamma}_{p}^{\prime}$, related to fermionic operators $\hat{f}_{p}$ and $\hat{f}_{p}^{\dagger}$ for $\psi$ fermions), via a Jordan-Wigner mapping (c).
$\hat{H}(t)=$ $\sum_{p}\Big{\{}\frac{g}{2}\left[1+(-1)^{\hat{b}_{v(p)}^{\dagger}\hat{b}_{v(p)}}\right]i\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime}$
$-d_{x}(t)i\hat{\gamma}_{p}\hat{\gamma}_{p+x}^{\prime}-d_{y}(t)\hat{\mathcal{S}}_{p}^{(y)}\ i\hat{\gamma}_{p}\hat{\gamma}_{p+y}^{\prime}\Big{\}}.$ (S20)
[…]
$\hat{H}_{\mathrm{tc}}^{\prime}=\frac{g}{2}(\hat{A}_{v}+\hat{A}_{v^{\prime}})-\frac{g}{2}\sum_{v^{\prime\prime}\neq v,v^{\prime}}\hat{A}_{v^{\prime\prime}}\ -\frac{g}{2}\sum_{p}\hat{B}_{p},$ (S21)
such that the state containing the $e$ pair is the ground state with respect to the modified Hamiltonian. While this has no effect on the driving terms $\hat{X}_{i}\hat{Z}_{j}$ and their quasiparticle mapping, it changes the free toric-code part of Eqs. (S15) and (S16), which now maps to
$\hat{H}_{\rm tc}^{\prime}\leftrightarrow\frac{g}{2}\sum_{v^{\prime\prime}\neq v,v^{\prime}}\left[1+(-1)^{\hat{b}_{v^{\prime\prime}}^{\dagger}\hat{b}_{v^{\prime\prime}}}\right]i\gamma_{p(v^{\prime\prime})}\gamma_{p(v^{\prime\prime})}^{\prime}\\ \frac{g}{2}\sum_{v^{\prime\prime}\in\{v,v^{\prime}\}}\left[1-(-1)^{\hat{b}_{v(p)}^{\dagger}\hat{b}_{v(p)}}\right]i\gamma_{p}\gamma_{p}^{\prime},$ (S22)
such that its expectation value with respect to the boson state |$\ket{\vec{n}^{e}}$, containing the $e$-boson pair at vertices $v$ and $v^{\prime}$, gives
$\bra{\vec{n}^{e}}\hat{H}_{\rm tc}^{\prime}\ket{\vec{n}^{e}}=g\sum_{p}i\gamma_{p}\gamma_{p}^{\prime}\ .$ (S23)
This assumption is not important for the non-Abelian physics studied but it slightly simplifies the analysis of the fermion Hamiltonian, which then has a homogeneous chemical potential for all $p$ also in the presence of $e$ quasiparticles, thus directly mapping to the simplest common models of spinless superconductors [26].
Under this assumption and including the drive, the Hamiltonian (S20), for a fixed configuration of $e$ bosons defined by a boson Fock state |$\ket{\vec{n}^{e}}$, gives the bulk fermion Hamiltonian $\hat{H}_{\psi}(t)=\bra{\vec{n}^{e}}\hat{H}_{\rm tc}+\hat{H}_{d}(t)\ket{\vec{n}^{e}}$ reading as
$\hat{H}_{\psi}(t)=\sum_{p}\Big{\{}gi\hat{\gamma}_{p}\hat{\gamma}_{p}^{\prime}-d_{x}(t)i\hat{\gamma}_{p}\hat{\gamma}_{p+x}^{\prime}\\ -d_{y}(t)\hat{\mathcal{S}}_{p}^{(y)}\ i\hat{\gamma}_{p}\hat{\gamma}_{p+y}^{\prime}\Big{\}}.$ (S24)
@@
Re-expressing the Majorana operators in terms of complex fermions $\hat{f}_{p}$ and $\hat{f}_{p}$ via Eq. (S14), defining the fermion chemical potential $\mu_{\psi}=2g$ and neglecting constant energy shifts then gives the fermion Hamiltonian
$\hat{H}_{\psi}(t)=\mu_{\psi}\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}\\ -d_{x}(t)\sum_{p}(\hat{f}_{p}\hat{f}_{p+x}+\hat{f}_{p}^{\dagger}\hat{f}_{p+x}+\mathrm{H.c.})\\ -d_{y}(t)\sum_{p}\hat{\mathcal{S}}_{p}^{(y)}(\hat{f}_{p}\hat{f}_{p+y}+\hat{f}_{p}^{\dagger}\hat{f}_{p+y}+\mathrm{H.c.}).$ (S25)
corresponding to Eq. (3) of the main text.
@@
Consider now $e$-boson (vortex) movement. This is induced by the single-spin operator $\hat{Z}_{i}$. In the quasiparticle mapping, for spins in the bulk, this is mapped to [24]
$\hat{Z}_{v,R}\leftrightarrow(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+x}^{\dagger}+\hat{b}_{v+x}),$
$\hat{Z}_{v,T}\leftrightarrow\hat{\tilde{\mathcal{S}}}_{v}^{(y)}(\hat{b}_{v}^{\dagger}+\hat{b}_{v})(\hat{b}_{v+y}^{\dagger}+\hat{b}_{v+y}),$ (S26)
involving a string of fermion operators
@@ II Floquet engineering
II.1 Effective Hamiltonian
In this section, the high-frequency Floquet Hamiltonian of Eq. (5) is derived. The latter is produced by the driven fermion Hamiltonian $\hat{H}_{\psi}(t)=\hat{H}_{\psi}(t+T)$ of Eq. (3), which we report here,
$\hat{H}_{\psi}(t)=\sum_{p}\Big{[}\mu_{\psi}\hat{f}_{p}^{\dagger}\hat{f}_{p}\\ -\!\!\sum_{r\in\{x,y\}}d_{r}(t){\mathcal{S}}_{p}^{(r)}\big{(}\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{]}.$ (S28)
Following Floquet’s theorem, the evolution operator generated by a time-periodic Hamiltonian $\hat{H}(t)=\hat{H}(t+T)$, starting from an initial time $t_{0}$, can be decomposed according to [31, 33, 29, 30]
$\hat{U}(t,t_{0})=\hat{U}_{F}(t,t_{0})e^{-i\hat{H}_{F}^{[t_{0}]}(t-t_{0})}$ (S29)
in terms of a time-independent Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]}$ and a time-periodic micromotion operator $\hat{U}_{F}(t,t_{0})=\hat{U}_{F}(t_{0}+T,t_{0})=e^{\hat{K}(t,t_{0})}$, generated by the ‘kick’ operator $\hat{K}(t,t_{0})$. The stroboscopic dynamics in steps of $T$ is then fully ruled by $\hat{H}_{F}^{[t_{0}]}$ according to
$\hat{U}(t_{0}+nT,t_{0})=e^{-inT\hat{H}_{F}^{[t_{0}]}}.$ (S30)
The superscript in $\hat{H}_{F}^{[t_{0}]}$ makes explicit the parametric dependence on the initial time $t_{0}$ of the evolution and thus on the global phase of the drives at that time. As detailed in the following, for the driving scheme analyzed here, the stroboscopic generators at different $t_{0}$ (as well as the non-stroboscopic Floquet Hamiltonian [34]) match to leading order in high frequency and exhibit the desired topological properties. For this reason, in the main text we dropped the superscript on the Floquet Hamiltonian and denoted it with $\hat{H}_{F}$ (the impact of the phase of the drives on the non-Abelian vortex exchange phase is quantitatively analyzed below in the next section). Both $\hat{H}_{F}^{[t_{0}]}$ and $\hat{K}(t,t_{0})$ can be determined order by order in a high-frequency expansion in powers of $\omega^{-1}$ [34, 29]. The leading order is given by
@@ II.1 Effective Hamiltonian
$\hat{K}(t,t_{0})=-\sum_{m\neq 0}\frac{\left(e^{im\omega t}-e^{im\omega t_{0}}\right)\hat{H}_{m}}{m\omega},$ (S31b)
where we have indicated the Fourier components of $\hat{H}(t)$ with
@@
To compute the expansion for the Floquet Hamiltonian produced by $\hat{H}_{\psi}(t)$ of Eq. (S28), it is convenient to first represent $\hat{H}_{\psi}(t)$ in a rotating frame defined by the gauge transformation
$\hat{R}(t)=\exp\left(-it\frac{\omega}{2}\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}\right),$ (S33)
such that the evolution operator decomposes as $\hat{U}(t,t_{0})=\hat{R}(t)\hat{V}(t,t_{0})$. At the end of one period, $\hat{R}(T)$ reduces to the total fermion parity of the system, $\hat{R}(T)=\exp\left(-i\pi\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}\right)=(-1)^{\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}}$, which is a conserved quantity at all times, $[\hat{R}(T),\hat{H}_{\psi}(t)]=0$. Moreover, $\hat{R}(T)$ reduces to the identity within the physical subspace of even-parity states, and thus $\hat{R}(t)$ is $T$-periodic in that subspace. The transformed Hamiltonian $\hat{H}^{\prime}_{\psi}(t)=\hat{R}^{\dagger}(t)\hat{H}_{\psi}(t)\hat{R}(t)-i\hat{R}^{\dagger}(t)\partial_{t}\hat{R}(t)$, generating $\hat{V}(t,t_{0})$, is also $T$-periodic and reads
@@
$-\!\!\sum_{r\in\{x,y\}}d_{r}(t)\hat{\mathcal{S}}_{p}^{(r)}\big{(}e^{i\omega t}\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{]}.$ (S34)
Applying Floquet’s theorem to $\hat{H}^{\prime}_{\psi}(t)$, the propagator $\hat{V}(t,t_{0})$ can be decomposed as in Eq. (S29) with a Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]\prime}$. Inserting the driving modulation $d_{r}(t)=J+2\Delta\cos(\omega t+\phi_{r})$ of Eq. (4) into Eq. (S34), the Hamiltonian $\hat{H}^{\prime}_{\psi}(t)$ possesses the following time dependence,
$\hat{H}^{\prime}_{\psi}(t)=\sum_{p}\Big{\{}(\mu_{\psi}-\omega/2)\hat{f}_{p}^{\dagger}\hat{f}_{p}-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}J\big{(}e^{i\omega t}\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\\ -\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\Delta\big{[}e^{i\phi_{r}}\big{(}\hat{f}_{p}\hat{f}_{p+r}+e^{i\omega t}(\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\hat{f}_{p+r}^{\dagger}\hat{f}_{p})+e^{i2\omega t}\hat{f}_{p}\hat{f}_{p+r}\big{)}+\mathrm{H.c.}\big{]}\Big{\}}.$ (S35)
The leading order of the high-frequency expansion for the Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]\prime}$ is obtained by time averaging [Eq. (S31a)], which yields the time-independent part of Eq. (S35), namely
$\hat{H}_{F}^{[t_{0}]\prime}=\sum_{p}\Big{\{}-\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}-\!\!\sum_{r\in\{x,y\}}\hat{\mathcal{S}}_{p}^{(r)}\big{(}J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}\\ +\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+\mathrm{H.c.}\big{)}\Big{\}},$ (S36)
where we have defined the effective chemical potential $\mu=\omega/2-\mu_{\psi}$. In the original reference frame [before transforming through $\hat{R}(T)$], the Floquet Hamiltonian $\hat{H}_{F}^{[t_{0}]}$ is determined from the relation
$e^{-i\hat{H}_{F}^{[t_{0}]}T}$ $=\hat{U}(t_{0}+T,t_{0})=\hat{R}(T)\hat{V}(t_{0}+T,t_{0})$
$=e^{-i\pi\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}}e^{-i\hat{H}_{F}^{[t_{0}]\prime}T}$ (S37)
and the fact that $[\hat{R}(T),\hat{H}_{F}^{[t_{0}]\prime}]=0$ due to the conservation of the total fermion parity. The effect of $\hat{R}(T)$ is thus simply to shift the zero Bogoliubov–de Gennes quasienergy by $\pi/T$, that is half a Floquet-Brillouin zone. By choosing the gauge $\hat{f}_{p}\to e^{-i\phi_{x}/2}\hat{f}_{p}$, it becomes clear that the effective Hamiltonian of Eq. (S36) only depends, to leading order in high-frequency, on the phase-delay $\phi_{y}-\phi_{x}$ between the horizontal and vertical drives, and not on the global phase of the drives (the impact of the latter is analyzed in the following section). The kick operator $\hat{K}(t,t_{0})$, to leading order, is determined by Eq. (S31b), given $\hat{R}(t)$ and the Fourier components of $\hat{H}^{\prime}_{\psi}(t)$. The latter, from Eq. (S35), read